Sol Summary – February 14th

Sol: 4

Summary Title: Nothing Says Love Like Bacon!

Author’s name: Dave Masaitis

Mission Status: Nominal

Sol Activity Summary: Crew 205 awoke to overcast skies and the smell of bacon and eggs, as the Commander and Executive Officer cooked a breakfast of bacon, eggs, and toast prepared from homemade bread. Unfortunately, the only sunny things today were the egg yolks. The EVA team for EVA #07 left on schedule, but was recalled 45 minutes later as weather began to turn inclement. The EVA crew was in the airlock before any unsafe ground conditions occured, but they only completed the UAV test flight, and did not make it to the North Ridge. The rest of the crew helped the EVA team clean and store their EVA suits, before everyone headed upstairs for lunch. While having lunch, the crew engaged in a "Secret Valentine" gift exchange and then used the afternoon to catch up on scientific analysis of samples collected over the week’s EVAs so far. Towards the evening, the crew gathered in the upstairs common area of the Hab to prepare reports for the communications window, share stories, and discuss what would be prepared for dinner. After the communications window, the crew will eat chicken and rice casserole, discuss tomorrow’s priorities of work, and then play some card games before bed.

Look Ahead Plan: Due to today’s ongoing inclement weather, the ground tomorrow will be too saturated to make EVAs practical. Instead, the crew will remain at the Hab to focus on refining research, improving ongoing investigations, and drafting procedural documents to improve the workflow of future crews. The inability to conduct EVAs can be frustrating, but having multiple crew members with culinary skills makes being stuck at the Hab much more palatible. Tomorrow will also see another rotation of Commander and Executive Officer, so a new day will bring new leadership.

Anomalies in work: EVA# 07 terminated early due to weather

Weather: Cloudy all day long, rain started after 1100 hrs, with consistent rain all afternoon and into the evening

Crew Physical Status: Nominal


EVA#07: 0518800/4250900 (Pooh’s Corner), 0518600/4251500 (North Ridge, not reached)

Reports to be filed: Sol Summary, EVA report, Operations Report, Greenhab Report, Journalist Report

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site